# SETA: Semantic-Aware Token Augmentation for Domain Generalization

### v1

@@ V Conclusion
In the paper, we propose a token-level augmentation method SETA for DG to induce shape bias of the models. Different from prior CNN-based works that focus on diversifying style information within the channel dimension, our method involves discerning edge tokens and perturbing local edge features within the spatial dimension. This strategy compels the model to exhibit increased sensitivity to shape cues and improves its generalizability. To further boost model robustness to domain shifts, we extend the SETA to two advanced stylized variants with SOTA augmentation methods. Comprehensive experiments demonstrate the effectiveness of our method across various public datasets on different ViT and MLP architectures. In future work, we will further investigate the shape bias of DG models and develop advanced adaptive augmentation methods to enhance the shape bias and generalization ability of different advanced architectures.

### v2

@@ V Conclusion
In the paper, we propose a token-level augmentation method SETA for DG to induce shape bias of the models. Different from prior CNN-based works that focus on diversifying style information within the channel dimension, our method involves discerning edge tokens and perturbing local edge features within the spatial dimension. This strategy compels the model to exhibit increased sensitivity to shape cues and improves its generalizability. To further boost model robustness to domain shifts, we extend the SETA to two advanced stylized variants with SOTA augmentation methods. Comprehensive experiments demonstrate the effectiveness of our method across various public datasets on different ViT and MLP architectures. However, since SETA uses a Fourier-based low-frequency filtering method, its effectiveness for edge extraction might be influenced by the complexity of image textures, which could affect the distinction between class-relevant edges and noise. To address this, exploring adaptive methods for feature division and extraction is a promising direction to enhance the model’s shape bias further. In future work, we will explore advanced methods that can learn to adaptively refine shape-irrelated edge features from complex textures to enhance generalization. SETA could also be seamlessly integrated into image encoders of ViT-based vision-language models, which helps the network focus on domain-invariant shape semantics. We will also further improve our SETA by leveraging text information to align visual and textual representations and help the model better learn generalizable semantic information.